# Weaker quantization dimension results for self-similar measures

### v1

@@ Abstract.
In this paper, we investigate the quantization dimension of self-similar measures, particularly when the IFS does not satisfy the separation condition, but the sub-IFS at some level satisfies the separation condition. Further, we study the approximation of the space of Borel probability measures $\mathcal{P}(\mathbb{R}^{m})$ with respect to the geometric mean error, i.e., the quantization dimension of order zero.
@@
The first author is supported by the SEED Grant Project of IIIT Allahabad, India. The second and third authors are financially supported by the Ministry of Education, India.
@@ 1. Introduction
An important factor in the study of fractal geometry is the fractal dimension. It measures a fractal’s scaling across various magnifications to provide information on geometric complexity; it translates irregularity into an exact numerical quantity, the fractal dimension, that interpolates between traditional geometric dimensions. Since the irregularity exhibited by the fractal is not always homogeneous, a single notion of the fractal dimension is not sufficient. There are various dimensions that have been introduced in the literature to capture the different aspects of the complexity, including the Hausdorff dimension $\dim_{\mathcal{H}}(\cdot),$ Minkowski–Bouligand dimension (box-counting dimension) $\dim_{B}(\cdot),$ packing dimension $\dim_{P}(\cdot),$ Assouad dimension $\dim_{A}(\cdot),$ lower Assouad dimension (lower dimension) $\dim_{L}(\cdot),$ quantization dimension $\dim_{r}(\cdot),~L^{q}$- dimension, information dimension, correlation dimension, entropy dimension, etc. (see [4, Fraser, 12, Mat3, Sh1, 24, 18]). Some dimensions, such as the Hausdorff and box dimensions, capture global scaling complexity, while others, such as the Assouad and lower dimensions, describe local extremal behaviour. In the case of fractal measures, global notions such as the information dimension and the $L^{q}$-dimension, as well as local dimensions such as the local dimension and the multifractal spectrum, provide insight at the local level. The concept of quantization was first derived in information theory and signal processing, where it refers to the approximation of a continuous signal by a finite set of discrete code points. Zadore [24] traces this idea and derived the asymptotic formulas for the optimal quantization error at high rates of a probability distribution. In this context, a natural question is to determine the rate of convergence of the optimal-$n$ approximation error as $n$ approaches infinity. Building on these ideas, Graf and Luschgy [6] developed a measure-theoretic framework for quantization of probability distribution and formalized the notion of quantization coefficient and quantization dimension. In general, calculating the fractal dimension exactly is a tedious task. It becomes easier when the cylinders have no or minimal overlap. There are various separation properties in the literature, such as strong separation condition (SSC), open set condition (OSC), strong open set condition (SOSC), and exponential separation condition (ESC).
@@ 2. Preliminaries
(see [Mat, 17]). If $(X,d)$ is a compact metric space, then $(\Omega(X),d_{L})$ is a compact metric space (see [2, Theorem 5.1]).
@@ 3. Main Results
Let $\mu$ be a self-similar measure associated with a WIFS $(\{f_{i}\}_{i=1}^{N};\rho_{1}\\ ,\rho_{2},\ldots,\rho_{N})$ and $\mu_{n}$ be the self-similar measure associated with sub-WIFS $\mathcal{I}_{n}$ at $n$-th level. Then, $\mu_{n}$ is absolutely continuous with respect to $\mu$ i.e., $\mu_{n}<<\mu.$
@@
A direct comparison of Equations 3.1 and 3.2 shows that Equation 3.1 can be derived from Equation 3.2 by omitting certain terms in its summation. Also, utilising the fact that the composition of similarity mapping is again a similarity mapping and $\mu$ is a Borel probability measure, we are able to establish our claim. Next, we claim that $\mathfrak{P}^{k}(\mu)$ converges setwise to $\mu_{n}$. Setting $\lambda=\mu$ and following the notation in [23, Theorem 1], the sequence $\{\mathfrak{P}^{k}(\mu)\}_{k\in\mathbb{N}}$ satisfies its hypotheses. Hence, $\mathfrak{P}^{k}(\mu)(A)\to\mu_{n}(A)$ for all Borel sets $A$, completing the proof.
@@
Let $\mu$ be a self-similar measure associated with a WIFS $(\{f_{i}\}_{i=1}^{N}\\ ;\rho_{1},\rho_{2},\ldots,\rho_{N})$ and $\mu_{n}$ be the sub-IFS at $n$-th level. Then, $\mu_{n}$ is absolutely continuous with respect to $\mu,$ i.e., $\mu_{n}<<\mu$ in $L_{\infty}$ density.
@@
In order to prove the absolute continuity of measure $\mu_{n}$ with respect to measure $\mu$ in $L_{\infty}$ density, it is sufficient to show that $\mu_{n}(A)\leq M\mu(A)$ for some fixed constant $M>0$ and for all Borel measurable sets $A$. By Lemma 3.1 the sequence $\mathfrak{P}^{k}(\mu)$ converges setwise to $\mu_{n}$ and $\mathfrak{P}^{k}(\mu)\leq\mu$ for all $k\in\mathbb{N}$. This implies that
@@
In Theorem 3.1 and Corollary 3.2, we have shown that $\mu_{n}$ is absolutely continuous with respect to $\mu$ in $L_{\infty}$ density. Hence, by [22, Theorem 3.4], we obtained $D_{r}(\mu_{n})\leq D_{r}(\mu).$
@@
By [6, Corollary 11.4], for $r\in[1,\infty),$ the quantization dimension function is a non-decreasing function in $r$. One can observe that this result holds for $r\in[0,\infty)$ due to the fact that for $0\leq p\leq q<\infty,~~\|x\|_{p}\leq\|x\|_{q}.$
@@
Let $\mu$ be a self similar measure associated with an WIFS $(\{f_{i}\}_{i=1}^{N};\\ \{\rho_{i}\}_{i=1}^{N}).$ Then, by [7, Lemma 3.5]
@@
This depicts that $D_{0}(\mu+\nu)\geq\max\{D_{0}(\mu),D_{0}(\nu)\}.$ Now, let $t>\max\{D_{0}(\mu),D_{0}(\nu)\},$ then by [7, Proposition 4.3], we have
@@
Following a similar approach to [22, Lemma 3.10], we obtain the required result. ∎
Using the above lemmas, we will approximate the set $\mathcal{P}(\mathbb{R}^{m})$ by several of its subclasses with respect to the quantization dimension.
@@
The sets $\mathcal{P}_{0}({\mathbb{R}^{m}})=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\lim_{r\to 0^{+}}D_{r}(\mu)=D_{0}(\mu)=\dim_{H}(\mu)\}$ and $\mathcal{P}_{0>}({\mathbb{R}^{m}})=\{\mu\in\mathcal{P}(\mathbb{R}^{m}):\lim_{r\to 0^{+}}D_{r}(\mu)=D_{0}(\mu)>\dim_{H}(\mu)\}$ are dense in $\mathcal{P}(\mathbb{R}^{m}).$
@@
Since By [7, Remark 4.6], the set $\mathcal{P}_{0}({\mathbb{R}^{m}})$ is non-empty. Let $\mu\in\mathcal{P}(\mathbb{R}^{m})$ be any Borel probability measure and $\epsilon>0,$ then by the fact that $\mathcal{L}(\mathbb{R}^{m})$ is dense in $\mathcal{P}(\mathbb{R}^{m}),$ there exists a measure $\nu\in\mathcal{L}(\mathbb{R}^{m})$ such that
@@
Further, by Lemma 3.10 and the similar technique followed in [22, Theorem 3.16], we have there exists a compactly supported measure such that $d_{L}(\nu,\lambda)<\frac{\epsilon}{2},$ where $\lambda=\nu*\lambda_{1}$ for some $\lambda_{1}\in\mathcal{P}_{0}({\mathbb{R}^{m}}).$ in this way, we have
@@
This proves that the set $\mathcal{P}_{0}({\mathbb{R}^{m}})$ is a dense subset of $\mathcal{P}(\mathbb{R}^{m}).$ Again, by theorem [7, Theorem 4.3]the set $\mathcal{P}_{0>}({\mathbb{R}^{m}})$ is non-empty. Since the proof is similar, we omit it.
3.1. Approximation of measures in terms of quantization dimension
We introduce the total variation metric ([17]) on $\mathcal{P}(\mathbb{R}^{m})$ as follows:
@@ 3.1. Approximation of measures in terms of quantization dimension
Let $\lambda\in\mathcal{P}(\mathbb{R}^{m})$ be a fixed measure and $\big(\mu_{n}\big)$ be a sequence of absolutely continuous (with respect to $\lambda$) measures in $\mathcal{P}(\mathbb{R}^{m}).$ If $(\frac{d\mu_{n}}{d\lambda})$ converges in $\mathcal{L}^{1}(\lambda)$-norm then $\big(\mu_{n}\big)$ converges to a measure $\mu$, and
@@
is in $\mathcal{P}(\mathbb{R}^{m}).$ Further, we obtain
@@
Let $\lambda\in\mathcal{P}(\mathbb{R}^{m}).$ Suppose $\mu$ is a absolutely continuous measure with respect to $\lambda$ such that $\dim_{H}\Big(Gr\big(\frac{d\mu}{d\lambda}\big)\Big)=\alpha$ for some $m\leq\alpha\leq m+1.$ Then there exists a sequence of absolutely continuous measures $(\mu_{n})\in\mathcal{P}(\mathbb{R}^{m})$ with the following conditions:
@@
In the light of an analogous result to Theorem LABEL:densethm, we may have a sequence of integrable (with respect to $\lambda$) functions $(g_{n})$ such that $\dim_{H}\big(Gr(g_{n})\big)=\alpha$ and $g_{n}\to\frac{d\mu}{d\lambda}$ in $\mathcal{L}^{1}(\lambda)$-norm. Now, one defines a sequence of measures $(\mu_{n})$ as follows

### v2

@@ Abstract.
In this paper, we investigate the quantization dimension of self-similar measures, particularly when the IFS does not satisfy the separation condition, but the sub-IFS at some level satisfies the separation condition. Further, we study the approximation of the space of Borel probability measures $\Omega(\mathbb{R}^{m})$ with respect to the geometric mean error, i.e., the quantization dimension of order zero.
@@
The first author is supported by the SEED Grant Project of IIIT Allahabad, India. The second author is financially supported by the Ministry of Education, India.
@@ 1. Introduction
An important factor in the study of fractal geometry is the fractal dimension. It measures a fractal’s scaling across various magnifications to provide information on geometric complexity; it translates irregularity into an exact numerical quantity, the fractal dimension, that interpolates between traditional geometric dimensions. Since the irregularity exhibited by the fractal is not always homogeneous, a single notion of the fractal dimension is not sufficient. There are various dimensions that have been introduced in the literature to capture the different aspects of the complexity, including the Hausdorff dimension $\dim_{\mathcal{H}}(\cdot),$ Minkowski–Bouligand dimension (box-counting dimension) $\dim_{B}(\cdot),$ packing dimension $\dim_{P}(\cdot),$ Assouad dimension $\dim_{A}(\cdot),$ lower Assouad dimension (lower dimension) $\dim_{L}(\cdot),$ quantization dimension $\dim_{r}(\cdot),~L^{q}$- dimension, information dimension, correlation dimension, entropy dimension, etc. (see [4, 7, 14, 20, 25, 29, 22]). Some dimensions, such as the Hausdorff and box dimensions, capture global scaling complexity, while others, such as the Assouad and lower dimensions, describe local extremal behaviour. In the case of fractal measures, global notions such as the information dimension and the $L^{q}$-dimension, as well as local dimensions such as the local dimension and the multifractal spectrum, provide insight at the local level. The concept of quantization was first derived in information theory and signal processing, where it refers to the approximation of a continuous signal by a finite set of discrete code points. Zadore [29] traces this idea and derived the asymptotic formulas for the optimal quantization error at high rates of a probability distribution. In this context, a natural question is to determine the rate of convergence of the optimal-$n$ approximation error as $n$ approaches infinity. Building on these ideas, Graf and Luschgy [8] developed a measure-theoretic framework for quantization of probability distribution and formalized the notion of quantization coefficient and quantization dimension. In general, calculating the fractal dimension exactly is a tedious task. It becomes easier when the cylinders have no or minimal overlap. There are various separation properties in the literature, such as strong separation condition (SSC), open set condition (OSC), strong open set condition (SOSC), and exponential separation condition (ESC).
@@ 2. Preliminaries
(see [19, 21]). If $(X,d)$ is a compact metric space, then $(\Omega(X),d_{L})$ is a compact metric space (see [2, Theorem 5.1]), also note that if $X$ is separable and complete then the space $\omega(\mathbb{R}^{m})$ is complete.
@@ 3. Main Results
Let $\mu$ be a self-similar measure associated with a WIFS $(\{f_{i}\}_{i=1}^{N};\\ \rho_{1},\rho_{2},\ldots,\rho_{N})$ and $\mu_{n}$ be the self-similar measure associated with sub-WIFS $\mathcal{I}_{n}$ at $n$-th level. Then, $\mu_{n}$ is absolutely continuous with respect to $\mu$ i.e., $\mu_{n}<<\mu.$
@@
A direct comparison of Equations 3.1 and 3.2 shows that Equation 3.1 can be derived from Equation 3.2 by omitting certain terms in its summation. Also, utilising the fact that the composition of similarity mapping is again a similarity mapping and $\mu$ is a Borel probability measure, we are able to establish our claim. Next, we claim that $\mathfrak{P}^{k}(\mu)$ converges setwise to $\mu_{n}$. Setting $\lambda=\mu$ and following the notation in [28, Theorem 1], the sequence $\{\mathfrak{P}^{k}(\mu)\}_{k\in\mathbb{N}}$ satisfies its hypotheses. Hence, $\mathfrak{P}^{k}(\mu)(A)\to\mu_{n}(A)$ for all Borel sets $A$, completing the proof.
@@
Let $\mu$ be a self-similar measure associated with a WIFS $(\{f_{i}\}_{i=1}^{N}\\ ;\rho_{1},\rho_{2},\ldots,\rho_{N})$ and $\mu_{n}$ be the sub-IFS at $n$-th level. Then, $\mu_{n}$ is absolutely continuous with respect to $\mu,$ i.e., $\mu_{n}<<\mu$ with an $L_{\infty}$ density.
@@
In order to prove the absolute continuity of measure $\mu_{n}$ with respect to measure $\mu$ with an $L_{\infty}$ density, it is sufficient to show that $\mu_{n}(A)\leq M\mu(A)$ for some fixed constant $M>0$ and for all Borel measurable sets $A$. By Lemma 3.1 the sequence $\mathfrak{P}^{k}(\mu)$ converges setwise to $\mu_{n}$ and $\mathfrak{P}^{k}(\mu)\leq\mu$ for all $k\in\mathbb{N}$. This implies that
@@
In Theorem 3.1 and Corollary 3.2, we have shown that $\mu_{n}$ is absolutely continuous with respect to $\mu$ with an $L_{\infty}$ density. Hence, by [27, Theorem 3.4], we obtain $D_{r}(\mu_{n})\leq D_{r}(\mu).$
@@
By [8, Corollary 11.4], for $r\in[1,\infty),$ the quantization dimension function is a non-decreasing function in $r$. One can observe that this result holds for $r\in[0,\infty)$ due to the fact that for $0\leq p\leq q<\infty,~~\|x\|_{p}\leq\|x\|_{q}.$
@@
Let $\mu$ be a self similar measure associated with an WIFS $(\{f_{i}\}_{i=1}^{N};\\ \{\rho_{i}\}_{i=1}^{N}).$ Then, by [9, Lemma 3.5]
@@ 3.1. Approximation of measures in terms of quantization dimension
This depicts that $D_{0}(\mu+\nu)\geq\max\{D_{0}(\mu),D_{0}(\nu)\}.$ Now, let $t>\max\{D_{0}(\mu),D_{0}(\nu)\},$ then by [9, Proposition 4.3], we have
@@
Following a similar approach to [27, Lemma 3.10], we obtain the required result. ∎
Using the above lemmas, we will approximate the set $\Omega(\mathbb{R}^{m})$ by several of its subclasses with respect to the quantization dimension.
@@
The sets $\Omega_{0}({\mathbb{R}^{m}})=\{\mu\in\Omega(\mathbb{R}^{m}):\lim_{r\to 0^{+}}D_{r}(\mu)=D_{0}(\mu)=\dim_{H}(\mu)\}$ and $\Omega_{0>}({\mathbb{R}^{m}})=\{\mu\in\Omega(\mathbb{R}^{m}):\lim_{r\to 0^{+}}D_{r}(\mu)=D_{0}(\mu)>\dim_{H}(\mu)\}$ are dense in $\Omega(\mathbb{R}^{m}).$
@@
Since By [9, Remark 4.6], the set $\Omega_{0}({\mathbb{R}^{m}})$ is non-empty. Let $\mu\in\Omega(\mathbb{R}^{m})$ be any Borel probability measure and $\epsilon>0,$ then by the fact that $\mathcal{L}(\mathbb{R}^{m})$ is dense in $\Omega(\mathbb{R}^{m}),$ there exists a measure $\nu\in\mathcal{L}(\mathbb{R}^{m})$ such that
@@
Further, by Lemma 3.10 and the similar technique followed in [27, Theorem 3.16], we have there exists a compactly supported measure such that $d_{L}(\nu,\lambda)<\frac{\epsilon}{2},$ where $\lambda=\nu*\lambda_{1}$ for some $\lambda_{1}\in\Omega_{0}({\mathbb{R}^{m}}).$ in this way, we have
@@
This proves that the set $\Omega_{0}({\mathbb{R}^{m}})$ is a dense subset of $\Omega(\mathbb{R}^{m}).$ Again, by [9, Theorem 4.3] the set $\Omega_{0>}({\mathbb{R}^{m}})$ is non-empty. Since the proof is similar, we omit it.
We introduce the total variation metric ([21]) on $\Omega(\mathbb{R}^{m})$ as follows:
@@
Let $\lambda\in\Omega(\mathbb{R}^{m})$ be a fixed measure and $\big(\mu_{n}\big)$ be a sequence of absolutely continuous (with respect to $\lambda$) measures in $\Omega(\mathbb{R}^{m}).$ If $(\frac{d\mu_{n}}{d\lambda})$ converges in $\mathcal{L}^{1}(\lambda)$-norm then $\big(\mu_{n}\big)$ converges to a measure $\mu$, and
@@
is in $\Omega(\mathbb{R}^{m}).$ Further, we obtain
@@
Let $\lambda\in\Omega(\mathbb{R}^{m}).$ Suppose $\mu$ is a absolutely continuous measure with respect to $\lambda$ such that $\dim_{H}\Big(Gr\big(\frac{d\mu}{d\lambda}\big)\Big)=\alpha$ for some $m\leq\alpha\leq m+1.$ Then there exists a sequence of absolutely continuous measures $(\mu_{n})\in\Omega(\mathbb{R}^{m})$ with the following conditions:
@@
In the light of an analogous result to Theorem 3.12, we may have a sequence of integrable (with respect to $\lambda$) functions $(g_{n})$ such that $\dim_{H}\big(Gr(g_{n})\big)=\alpha$ and $g_{n}\to\frac{d\mu}{d\lambda}$ in $\mathcal{L}^{1}(\lambda)$-norm. Now, one defines a sequence of measures $(\mu_{n})$ as follows